I have to solve the radical equation sq of x+8 = sq of 5x-4
Please simplify (as I am not good at this) and explain how you did it.
Thank you.
Donna

Answer by dabanfield(803)

I have to solve the radical equation sq of x+8 = sq of 5x-4
Please simplify (as I am not good at this) and explain how you did it.
Thank you.
Donna
Hi Donna,
Is the problem sqrt(x+8) = sqrt(5x-4) or (x+8)^2 = (5x-4)^2 ?
Since you said "radical" I'm thinking it's
1.) sqrt(x+8) = sqrt(5x-4)
We know if a = b then a^2 = b^2 so we can square both sides of 1.) above:
2.) [sqrt(x+8)]^2 = [sqrt(5x-4)]^2
We also know that [sqrt(a)]^2 = a so 2.) becomes:
x+8 = 5x-4
Now you can solve the above for x.